what is the simplified form of (5/6)/(2/3)
Answer by fcabanski(1391)

(5/6)/(2/3) = 5/6 * 3/2 = 15/12 = 3/4

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)